The 2006th digit in sequence 123456789101112131415..........is

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
All the natural numbers starting from 1 are enlisted:
So number of 1 digit numbers = 9
number of 2 digit numbers = 90(to calculate number of digits we need to multiply by 2)
number of 3 digit numbers starting from 100 till 699 = 600(to calculate number of digits we need to multiply by 3)
So, starting from 1 to 699, total number of digits will be
 9 + 2*90 + 3*600 = 1989.
If we also add 700-704 we get total 1989 + 15 = 2004 digits enlisted.
Now the 2006 digit in the sequence is clearly evident that would be the second digit of 705 which is 0.
Regards
Arun (askIITians forum expert)